OpenStudy (anonymous):

A sphere of radius is illuminated by a light source that is 4m away from the sphere's center. What is the surface area of the sphere that is brightened?

OpenStudy (alekos):

whats the radius?

OpenStudy (anonymous):

oh sorry i forgot , it's 2m :)

OpenStudy (anonymous):

@alekos

OpenStudy (anonymous):

@Michele_Laino can you help me with this question :)

OpenStudy (alekos):

have you been given an answer?

OpenStudy (anonymous):

no but theses are the choices : a. pi b. 2 pi c 4 pi d 6pi

OpenStudy (michele_laino):

I think that your problem can be represented like below: 

OpenStudy (anonymous):

O.o

OpenStudy (michele_laino):

we have \(PQ=4\), \(OQ=2\), Then, using the 2-nd theorem of Euclid, we get \(OH=1\)

OpenStudy (michele_laino):

oops..\(OP=4\)

OpenStudy (michele_laino):

 then \(OS=2-1=1\) and hence the requested surface is: \(2 \cdot \pi \; \cdot OQ \;HS=2 \;\cdot \pi\;\cdot 2\;\cdot 1=4\;\pi\)

OpenStudy (anonymous):

thank you thank you thank you ^_^

OpenStudy (michele_laino):

:)

OpenStudy (michele_laino):

oops.. I meant \(HS=OS-OH=2-1=1\)